OpenStudy (anonymous):

Factor completely 25x2- 144 Would this be the answer below? (5x - 12)(5x - 12)

OpenStudy (anonymous):

No, not quite, the answer is (5x - 12)(5x + 12)

OpenStudy (anonymous):

wait: (5x-12)(5x+12) ?

OpenStudy (anonymous):

Hahaha, yeah. This is an identity you will see show up over and over: a^2 - b^2 = (a + b)(a - b)

OpenStudy (anonymous):

It's convenient because the "middle" (power of 1) factors cancel out. Distributing: (a + b)(a - b) = a^2 + (-ab) + ab + b^2 = a^2 + b^2

OpenStudy (anonymous):

I understand what your saying. Just dont know how to respond. :P thank you

OpenStudy (anonymous):

Yup, you'll see it over and over- so remember it! Makes things easier